OpenStudy (anonymous):

Pamela's age is two times Jiri's age. The sum of their ages is 51. What is Jiri's age?

OpenStudy (cyber405):

Pamela's age = X Jiri's age = Y X=2Y and Y+2Y =51 Now Y=... got it?

OpenStudy (anonymous):

51=P+1/2P 51/1.5=34 34= pamela's age 34/2=17 17 = jiri's age 17+34=51

OpenStudy (anonymous):

Thanks :)

OpenStudy (cyber405):

Welcome:)